### v1

@@ 4.2 Kernel Locator
The kernel locator locates the file ranges (start and end file offset) of the used kernels in the ML shared library. It identifies a list of file ranges that need to be retained in the shared library. The difficulty in locating kernels in the device code is that there is no public specification for the structure of the device code. Finding the exact location of used kernels requires in-depth reverse engineering of the device code, which can be overfitted to a specific version of the CUDA toolkit and error-prone. In addition, because device-launching kernels are not detected by the kernel detector, the kernel locator must also handle these kernels appropriately to avoid mistakenly removing them.
@@
To address these issues, we propose an approach to approximately locate the used kernels in the device code. Instead of finding the exact location of kernels, we find the location of the cubins that contain the kernels. A cubin is a CUDA binary file that contains kernel code. If a kernel is launched by another kernel, the two kernels are compiled into the same cubin nvi (e). Based on this observation, we can deduce that if a cubin contains a host-launching kernel, it also contains all the kernels in the kernel call graph starting from the host-launching kernel, including those device-launching kernels. Therefore, retaining a cubin that contains a host-launching kernel also retains all the kernels in the kernel call graph starting from the host-launching kernel, including those device-launching kernels as well.
@@
To find whether a cubin contains a host-launching kernel, we use the cuobjdump tool  nvi (a) to extract a list of cubin files from the shared library. For each cubin, we extract the kernels included in it using cuobjdump. If a cubin contains a used host-launching kernel, the whole cubin needs to be retained. In doing so, we also retain the device-launching kernels in the cubin. The next step is to locate the cubin in the shared library file, i.e., find the file range occupied by the cubin. This involves understanding the structure of the device code in the shared library. As shown in Figure 4, the device code in a shared library is organized as a list of regions. Each region includes a region header and a list of elements. Each element includes an element header and a cubin. A cubin extracted by cuobjdump has an index starting from 1 in its file name. This index is equal to the index of the element containing the cubin in the shared library. Using this index, we can map the cubin to the corresponding element in the shared library. In doing so, we can locate the file range occupied by the cubin.
@@
To maintain the integrity of the shared library, the kernel locator retains or removes the whole element containing a cubin. Finally, the kernel locator uses the following criteria to determine whether to retain an element; The element header has a field called compute-capability, which shows the GPU architecture the cubin is compiled for. We find that only the elements that match the GPU architecture can be loaded into the GPU memory. Therefore, if an element matches the GPU architecture which the ML workload is running on and contains a cubin that has used host-launching kernels, then we retain the element.
@@
Figure 4 also shows an example of the relationship between kernels, cubins and elements in the shared library. The used kernel matmul, which is detected by the kernel detector, is contained in a cubin as shown in the figure. The kernel may launch other kernels, forming a kernel call graph. All the kernels in the kernel call graph are contained in the same cubin. The cubin is in turn contained in an element in the shared library. By retaining the whole element, we ensure that all the kernels in the kernel call graph are retained, including the device-launching kernels in the call graph.
@@
Compaction: The file ranges occupied by elements that meet the criteria are retained in the ML shared library, while the rest are removed. This process is handled by Negativa’s compaction phase. In this phase, the unused files are zeroed out. Negativa then maps the file offsets to their original memory addresses where the original shared library is loaded into memory to retain memory address validity. More details on the compaction process can be found in our supplementary material.
@@ 5.1 Overview of Bloat in ML Frameworks
We execute the workloads listed in Table 1 running Negativa-ML to debloat the shared libraries used by the workloads. After debloating, we re-run the workloads using the debloated shared libraries to ensure that the workloads run correctly. The removed code is considered bloat as they are not necessary for the workloads. We then compare the total file sizes of the shared libraries before and after debloating to assess file size reductions. Moreover, for host code, we compare both code size and the number of functions before and after debloating. For device code, we also analyze code size reductions and the number of elements removed. The results are presented in Table 2.
For the small model MobileNetV2, training with PyTorch involves 113 shared libraries, totaling 3762 MB. After debloating, the total file size decreases by 55%. Notably, the host code size is reduced by 68%, with 93% of functions removed. The device code size decreases by 75%, with 98% of elements removed. The device code also accounts for the majority of the total file size and file size reduction. Inference MobileNetV2 with PyTorch shows similar reductions.
For the same MobileNetV2 model, TensorFlow uses more shared libraries compared to PyTorch, with 253 libraries for training and 251 for inference The total file sizes and reductions are smaller than PyTorch. The host code in TensorFlow presents interesting results: it has a larger size and a greater number of functions than PyTorch, yet the reduction is less significant, indicating that TensorFlow uses more host code and functions than PyTorch. Given the fact that PyTorch can train(inference) the same model with less host code and functions, this suggests there may be unnecessary function calls within TensorFlow’s host code—functions that are used but do not contribute meaningfully to the target ML workload. The device code in TensorFlow also shows a significant reduction in size and elements, with over 99% of elements removed.
For the medium model, Transformer, both PyTorch and TensorFlow show similar results to MobileNetV2, indicating that different models do not significantly affect the bloat in the ML frameworks. For the large model, LLama2, the frameworks vLLM and Transformers were used for inference, and despite differences in frameworks, the reductions in file size, host code and device code remain significant.
The size of device code is considerably larger than that of host code for all ML frameworks. Device code also exhibits a higher reduction in both size and element count for all ML frameworks. In particular, the element count reduction in device code exceeds 97% across all workloads, underscoring that device code is significantly more bloated than host code.
Table 2: Total file size, host code, device code and their reductions of all shared libraries in each ML framework. The table shows the original value of a metric and the reduction in percentage of the metric in parentheses. K=1,000.
Model Framework Operation #Lib. Total File Size/MB Host Code Device Code
@@
Summary. All ML frameworks show substantial reductions in both host code ($\geq{46\%}$) and device code ($\geq{66\%}$). Device code is notably more bloated than host code, contributing to the majority of the bloat within the ML frameworks.
@@ 5.2 Shared Library Level Analysis
In this section, we delve deeper into the distribution of bloat within host and device code at the shared library level. For each shared library used in the workloads, we calculate the host code size reduction and function count reduction in percentage. Similarly, for each shared library, we calculate the reduction percentage in device code size and element count as well. If a shared library does not have device code, we exclude it from the device code analysis.
Figure 5(a) illustrates the distribution of size reduction in host and device code. The distribution patterns of host and device code show distinct differences. The median reduction for host code size is approximately 25%, with many shared libraries showing a reduction from 0% to 10%. In contrast, the median reduction for device code size is almost 80%, significantly higher than that for host code, and its distribution is also more concentrated. Figure 5(b) depicts the distribution of reductions in function and element counts. Notably, all shared libraries exhibit an element reduction of over 80%. The concentrated distributions for device code size and element count reductions highlight that device code is considerably more bloated than host code in all shared libraries.
(a) Distribution of host code size reduction and device code size reduction.
(b) Distribution of host function count reduction and device element count reduction.
For each workload, we sorted the shared libraries by their absolute file size reductions in descending order. We found that the top 10% shared libraries contribute over 90% of the total size reduction for all ML frameworks. Reductions in host code size and device code size also follow a similar pattern. For instance, as illustrated in Figure 6, the Pareto chart for the PyTorch Training MobileNetV2 workload shows that among 113 shared libraries, the top 8 libraries account for 90% of the total file size reduction. This result suggests that bloat follows a Power Law distribution, with a few shared libraries containing the majority of the bloat.
@@
Summary. 50% of shared libraries in ML frameworks have device code reductions more than 80%. 10% of shared libraries contribute over 90% of the total size reduction.
@@ 5.3 Function and Element Level Analysis
To deepen the analysis, Table 3 presents reductions for the largest shared library used in each workload. All the four ML frameworks running different models use the common two shared libraries, torch_cuda.so and tensorflow_cc.so. These two libraries, which provide the core functionalities of the ML frameworks, also exhibit significant reductions in file size, device code, and host code. For the torch_cuda.so, the reduction in file size, host code size, and device code size are 76%, 91% and 82%, respectively. The host code in tensorflow_cc.so also presents the interesting results as we have discussed in §5.1: it has a much larger host code size and a greater number of functions than torch_cuda.so, yet the reduction is smaller. But the reductions in host code for tensorflow_cc.so is as significant as torch_cuda.so.
@@
Model Framework Operation Lib. Name File Size/MB Host Code Device Code
@@
TensorFlow Train tensorflow_cc.so 965 (43) 300 (59) 670K (51) 298 (79) 1,637 (100)*
Inference tensorflow_cc.so 965 (43) 300 (61) 670K (52) 298 (79) 1,637 (100)*
@@
TensorFlow Train tensorflow_cc.so 965 (43) 300 (59) 670K (51) 298 (79) 1,637 (100)*
Inference tensorflow_cc.so 965 (41) 300 (59) 670K (52) 298 (73) 1,637 (94)
@@
We also examine the reasons for removed elements in device code. Negativa-ML removes an element due to one of the following two reasons: Reason I: The element does not match the GPU architecture;Reason II: The element matches the GPU architecture, but it does not have any used kernels. We identify the reason for each removed element.
Figure 7 shows that for all workloads, over 80% of removed elements are due to Reason I, i.e., the element does not match the GPU architecture. The cause of these unmatched elements is that the device code in these shared libraries are built to support various GPU architectures, leading to numerous unnecessary elements for workloads on a specific GPU. For example, in our experiment, we observed a single shared library in the PyTorch framework contained elements for 6 different GPU architectures. If the library is deployed on a specific GPU architecture, the elements for the five other architectures are unnecessary. This reveals a deep reason for bloat: software bloat can stem from hardware.
Figure 7: Reasons of removed elements in device code.
Summary. Different workloads used a lot of common functions in the core shared libraries. Most elements removed in device code are due to the mismatch of GPU architecture.
5.4 Runtime Performance Analysis
Running workloads with debloated libraries reduce memory usage (both host and GPU) and execution time. The best improvement is observed in the PyTorch Inference MobileNetV2 workload, with reductions of 74.6% in peak host memory, 69.6% in peak GPU memory, and 44.6% in execution time. Inference workloads generally show better improvement than training workloads. Across all workloads, average absolute reductions are 2501 MB for peak host memory, 443 MB for peak GPU memory, and 2.6 seconds for execution time. Notably, the absolute execution time reduction remains relatively constant, regardless of the actual execution duration. This improvement stems from reduced code size, which decreases the time required to load the code into memory. This execution time improvement is especially impactful for tasks sensitive to cold start latency, such as serverless ML applications.
@@ 5.4 Runtime Performance Analysis
Model Framework Operation Peak Host Memory/MB Peak GPU Memory/MB Execution Time/s
@@
Summary. Debloating shared libraries significantly reduces both host and GPU memory usage. The time required to load the code into memory is also reduced, leading to shorter execution times.
5.5 Evaluation on Distributed Inference
In this section, we present debloating results for additional ML workloads, focusing on two LLM inference frameworks, vLLM and Transformers. We select the top 9 LLMs from the Hugging Face Open LLM Leaderboard Beeching et al. (2023) and deploy them on both frameworks using distributed inference with 8×A100 40GB GPUs. This setup aims to evaluate whether debloating works with distributed inference under a different hardware setup. Results are detailed in Table 7 in the Appendix.

### v3

@@ 3.2 Kernel Locator
The kernel locator locates the file ranges (start and end file offset) of the used kernels in the ML shared library. It identifies a list of file ranges that need to be retained in the shared library. The difficulty in locating kernels in the GPU code is that there is no public specification for the structure of the GPU code. Finding the exact location of used kernels requires in-depth analysis of the GPU code, which can be overfitted to a specific version of the CUDA toolkit and error-prone. In addition, because GPU-launching kernels are not detected by the kernel detector, the kernel locator must also handle these kernels appropriately to avoid mistakenly removing them.
@@
To address these issues, we propose an approach to approximately locate the used kernels in the GPU code. Instead of finding the exact location of kernels, we find the location of the cubins that contain the kernels. A cubin is a CUDA binary file that contains kernel code. If a kernel is launched by another kernel, the two kernels are compiled into the same cubin nvi (e). Based on this observation, we can deduce that if a cubin contains a CPU-launching kernel, it also contains all the kernels in the kernel call graph starting from the CPU-launching kernel, including those GPU-launching kernels. Therefore, retaining a cubin that contains a CPU-launching kernel also retains all the kernels in the kernel call graph starting from the CPU-launching kernel, including those GPU-launching kernels as well.
@@
To find whether a cubin contains a CPU-launching kernel, we use the cuobjdump tool  nvi (a) to extract a list of cubin files from the shared library. For each cubin, we extract the kernels included in it using cuobjdump. If a cubin contains a used CPU-launching kernel, the whole cubin needs to be retained. In doing so, we also retain the GPU-launching kernels in the cubin. The next step is to locate the cubin in the shared library file, i.e., find the file range occupied by the cubin. This involves understanding the structure of the GPU code in the shared library. As shown in Figure 4, the GPU code in a shared library is organized as a list of regions. Each region includes a region header and a list of elements. Each element includes an element header and a cubin. A cubin extracted by cuobjdump has an index starting from one in its file name. This index is equal to the index of the element containing the cubin in the shared library. Using this index, we can map the cubin to the corresponding element in the shared library. In doing so, we can locate the file range occupied by the cubin.
@@
To maintain the integrity of the shared library, the kernel locator retains or removes the whole element containing a cubin. Finally, the kernel locator uses the following criteria to determine whether to retain an element; The element header has a field called compute-capability, which shows the GPU architecture the cubin is compiled for. We find that only the elements that match the GPU architecture can be loaded into the GPU memory. Therefore, if an element matches the GPU architecture which the ML workload is running on and contains a cubin that has used CPU-launching kernels, then we retain the element.
@@
Figure 4 also shows an example of the relationship between kernels, cubins and elements in the shared library. The used kernel matmul, which is detected by the kernel detector, is contained in a cubin as shown in the figure. The kernel may launch other kernels, forming a kernel call graph. All the kernels in the kernel call graph are contained in the same cubin. The cubin is in turn contained in an element in the shared library. By retaining the whole element, we ensure that all the kernels in the kernel call graph are retained, including the GPU-launching kernels in the call graph.
@@
Compaction: The file ranges occupied by elements that meet the criteria are retained in the ML shared library, while the rest are removed. This process is handled by Negativa’s compaction phase. In this phase, the unused files are zeroed out. Negativa then maps the file offsets to their original memory addresses where the original shared library is loaded into memory to retain memory address validity. More details on the compaction process can be found in  Zhang & Ali-Eldin (2025).
@@ 4.1 Overview of Bloat in ML Frameworks
Table 2: Total file size, CPU code, GPU code and their reductions of all shared libraries in each ML framework. The table shows the original value of a metric and the reduction in percentage of the metric in parentheses. K=1,000.
Model Framework Operation #Lib. Total File Size/MB CPU Code GPU Code
@@
We execute the workloads listed in Table 1 running Negativa-ML to debloat the shared libraries used by the workloads. After debloating, we re-run the workloads using the debloated shared libraries to verify the correctness of debloating. Specifically, we first compared outputs by the original and debloated workloads, confirming they are essentially identical. Then, we compared final metrics such as test loss, validation loss or generated text for the original and debloated workloads. We found that there are no differences between the output and final metrics, demonstrating that the debloating process does not affect the correctness of the workloads. We omit these comparisons due to space.
We then compare the total file sizes of the shared libraries before and after debloating to assess file size reductions. Moreover, for CPU code, we compare both code size and the number of functions before and after debloating. For GPU code, we also analyze code size reductions and the number of elements removed. The results are presented in Table 2.
For the small model MobileNetV2, training with PyTorch involves 113 shared libraries, totaling 3762 MB. After debloating, the total file size decreases by 55%. Notably, the CPU code size is reduced by 68%, with 93% of functions removed. The GPU code size decreases by 75%, with 98% of elements removed. The GPU code also accounts for the majority of the total file size and file size reduction. Inference MobileNetV2 with PyTorch shows similar reductions.
For the same MobileNetV2 model, TensorFlow uses more shared libraries compared to PyTorch, with 253 libraries for training and 251 for inference The total file sizes and reductions are smaller than PyTorch. The CPU code in TensorFlow presents interesting results: it has a larger size and a greater number of functions than PyTorch, yet the reduction is less significant, indicating that TensorFlow uses more CPU code and functions than PyTorch. Given the fact that PyTorch can train(inference) the same model with less CPU code and functions, this suggests there may be unnecessary function calls within TensorFlow’s CPU code—functions that are used but do not contribute meaningfully to the target ML workload. The GPU code in TensorFlow also shows a significant reduction in size and elements, with over 99% of elements removed.
For the medium model, Transformer, both PyTorch and TensorFlow show similar results to MobileNetV2, indicating that different models do not significantly affect the bloat in the ML frameworks. For the large model, LLama2, the frameworks vLLM and Transformers were used for inference, and despite differences in frameworks, the reductions in file size, CPU code and GPU code remain significant.
The size of GPU code is considerably larger than that of CPU code for all ML frameworks. GPU code also exhibits a higher reduction in both size and element count for all ML frameworks. In particular, the element count reduction in GPU code exceeds 97% across all workloads, underscoring that GPU code is significantly more bloated than CPU code.
Summary. All ML frameworks show substantial reductions in both CPU code ($\geq{46\%}$) and GPU code ($\geq{66\%}$). GPU code is notably more bloated than CPU code, contributing to the majority of the bloat within the ML frameworks.
@@ 4.2 Shared Library Level Analysis
In this section, we delve deeper into the distribution of bloat within CPU and GPU code at the shared library level. For each shared library used in the workloads, we calculate the CPU code size reduction and function count reduction in percentage. Similarly, for each shared library, we calculate the reduction percentage in GPU code size and element count as well. If a shared library does not have GPU code, we exclude it from the GPU code analysis.
Figure 5(a) illustrates the violin plots of distribution of size reduction in CPU and GPU code. The distribution patterns of CPU and GPU code show distinct differences. The median reduction for CPU code size is approximately 25%, with many shared libraries showing a reduction from 0% to 10%. In contrast, the median reduction for GPU code size is almost 80%, significantly higher than that for CPU code, and its distribution is also more concentrated. Figure 5(b) depicts the distribution of reductions in function and element counts. Notably, all shared libraries exhibit an element reduction of over 80%. The concentrated distributions for GPU code size and element count reductions highlight that GPU code is considerably more bloated than CPU code in all shared libraries.
(a) Distribution of CPU code size reduction and GPU code size reduction.
(b) Distribution of CPU function count reduction and GPU element count reduction.
Figure 5: Violin plots of distribution of CPU and GPU code reduction in size and function count.
For each workload, we sorted the shared libraries by their absolute file size reductions in descending order. We found that the top 10% shared libraries contribute over 90% of the total size reduction for all ML frameworks. Reductions in CPU code size and GPU code size also follow a similar pattern. For instance, as illustrated in Figure 6, the Pareto chart for the PyTorch Training MobileNetV2 workload shows that among 113 shared libraries, the top 8 libraries account for 90% of the total file size reduction. This result suggests that bloat follows a Power Law distribution, with a few shared libraries containing the majority of the bloat.
@@
Summary. 50% of shared libraries in ML frameworks have GPU code reductions more than 80%. 10% of shared libraries contribute over 90% of the total size reduction.
@@ 4.3 Function and Element Level Analysis
Model Framework Operation Lib. Name File Size/MB CPU Code GPU Code
@@
TensorFlow Train tf_cc.so2 965 (43) 300 (59) 670K (51) 298 (79) 1,637 (100)1
Inference tf_cc.so2 965 (43) 300 (61) 670K (52) 298 (79) 1,637 (100)1
@@
TensorFlow Train tf_cc.so2 965 (43) 300 (59) 670K (51) 298 (79) 1,637 (100)1
Inference tf_cc.so2 965 (41) 300 (59) 670K (52) 298 (73) 1,637 (94)
@@
Model Framework Operation Peak CPU Memory/MB Peak GPU Memory/MB Execution Time/s
@@
To deepen the analysis, Table 3 presents reductions for the largest shared library used in each workload. All the four ML frameworks while running the different models use one of two shared libraries, torch_cuda.so and tensorflow_cc.so. These two libraries, which provide the core functionalities of the ML frameworks, also exhibit significant reductions in file size, GPU code, and CPU code. For the torch_cuda.so, the reduction in file size, CPU code size, and GPU code size are 76%, 91% and 82%, respectively. The CPU code in tensorflow_cc.so also presents the interesting results as we have discussed in §4.1: it has a much larger CPU code size and a greater number of functions than torch_cuda.so, yet the reduction is smaller. The reductions in GPU code for tensorflow_cc.so is as significant as torch_cuda.so.
$J(A,B)=\frac{|A\cap B|}{|A\cup B|}$ (1)
Figure 7: Reasons of removed elements in GPU code.
We also examine the reasons for removed elements in GPU code. Negativa-ML removes an element due to one of the following two reasons: Reason I: The element does not match the GPU architecture; Reason II: The element matches the GPU architecture, but it does not have any used kernels. We identify the reason for each removed element.
Figure 7 shows that for all workloads, over 80% of removed elements are due to Reason I, i.e., the element does not match the GPU architecture. The cause of these unmatched elements is that the GPU code in these shared libraries are built to support various GPU architectures, leading to numerous unnecessary elements for workloads on a specific GPU. For example, in our experiment, we observed a single shared library in the PyTorch framework contained elements for 6 different GPU architectures. If the library is deployed on a specific GPU architecture, the elements for the five other architectures are unnecessary. This hints at a new reason for bloat: software bloat can stem from hardware.
Summary. Different workloads used a lot of common functions in the core shared libraries. Most elements removed in GPU code are due to the mismatch of GPU architecture.
4.4 Runtime Performance Analysis
Running workloads with debloated libraries reduces memory usage (both CPU and GPU) and execution time. The best improvement is observed in the PyTorch Inference MobileNetV2 workload, with reductions of 74.6% in peak CPU memory, 69.6% in peak GPU memory, and 44.6% in execution time. Inference workloads generally show better improvement than training workloads. Across all workloads, average absolute reductions are 2501 MB for peak CPU memory, 443 MB for peak GPU memory, and 2.6 seconds for execution time. Notably, the absolute execution time reduction remains relatively constant, regardless of the actual execution duration. This improvement stems from reduced code size, which decreases the time required to load the code into memory. This execution time improvement is especially impactful for tasks sensitive to cold start latency, such as serverless ML applications.
@@ 4.4 Runtime Performance Analysis
Summary. Debloating shared libraries significantly reduces both CPU and GPU memory usage. The time required to load the code into memory is also reduced, leading to shorter execution times.
4.5 Evaluation on Different GPUs
Table 6: Total file size, CPU code, GPU code and their reductions of all shared libraries in vLLM/Transformers inference Llama2 workload on 1 H100 GPU under eager-loading and lazy-loading modes. The table shows the original value of a metric and the reduction in percentage of the metric in parentheses. K=1000.
Framework Mode #Lib. Total File Size/MB CPU Code GPU Code
Size/MB #Functions Size/MB #Elements
vLLM Eager 170 4,068 (47) 726 (66) 872K (93) 2,104 (68) 7,692 (89)
Lazy 170 4,068 (47) 726 (66) 872K (93) 2,104 (68) 7,692 (89)
Transformers Eager 93 2,868 (51) 513 (72) 584K (92) 1,592 (69) 7,165 (88)
Lazy 93 2,868 (51) 513 (72) 584K (92) 1,592 (69) 7,165 (88)
Table 7: Average runtime performance using original shared libraries and reductions using debloated shared libraries in vLLM/Transformers inference Llama2 workload on 1 H100 GPU under eager-loading and lazy-loading modes. The numbers in parentheses are the percentage of reduction. The standard deviation of all metrics for each workload is less than 2% and is not shown.
Framework Mode Peak CPU Memory/MB Peak GPU Memory/MB Execution Time/s
vLLM Eager 13,333 (18.0) 91,946 (0.7) 44 (13.9)
Lazy 10,313 (0.3) 91,061 (0.0) 36 (8.3)
Transformers Eager 12,345 (12.2) 15,002 (2.4) 23 (32.0)
Lazy 10,049 (0.2) 13,764 (0.0) 17 (20.3)
In this section, we present debloating results of the two LLM inference frameworks, vLLM and Transformers, using different GPUs. We first evaluate the vLLM/Transformers inference workload using Llama2 on a single H100 GPU. The evaluation is conducted under both eager-loading and lazy-loading modes. In eager-loading, all kernels are loaded into memory at application startup, whereas lazy-loading loads kernels only when needed, reducing memory footprint. Table 6 presents the size reductions achieved. For the same framework, the reductions in file size, CPU code size, function count, and GPU code size are similar between the two loading modes. The runtime performance improvements are shown in Table 7. Under eager-loading, debloating results in greater CPU memory reduction compared to lazy-loading, while GPU memory reduction remains similar across both modes. Both loading modes exhibit reductions in execution time. Comparing these results with those obtained using the T4 GPU (Table 2 and Table 5), we observe consistent reductions across different GPUs. This demonstrates that Negativa-ML effectively debloats shared libraries across various GPU architectures.